OpenStudy (anonymous):

Use your calculator to find the approximate volume in cubic units of the solid created when the region under the curve y = sec(x) on the interval [0,pi divided by 3] is rotated around the x-axis. 1.317 1.732 4.137 5.441

OpenStudy (perl):

first i would graph y = pi * sec(x)^2 or y = pi /cos(x)^2 that will give you the cross sectional area at a point x. then just integrate from 0 to pi/3